# The Morse complex of the wedge of two extended star graphs and a path

### v1

@@ 2. Preliminaries
We write $V(K)\coloneqq\{\sigma^{(0)}\in K\}$ and refer to $V(K)$ as the vertex set of $K$. A 0-simplex adjacent to $v\in V(K)$ is called a neighbor of $v$. The degree of a vertex $v\in V(K)$, denoted $\deg(v)$ is the number of neighbors of $v$. If $\deg(v)=1$ and the neighbor of $v$ is $w$, we say that $\{v,vw\}$ is a leaf.
@@
By using Hasse diagrams, we can give an equivalent definition of the Morse complex as in Definition 2.6. We first define an acyclic matching.
Definition 2.7.
If a subset $M\subseteq P\times P$ satisfies the following two conditions for a poset $P$, then $M$ is called a partial matching.
  1. (i)
    For any $(a,b)\in M$, we have $a\prec b$, where $a\prec b$ means $a<b$ and there is no $c\in P$ such that $a<c<b$.
  2. (ii)
    Each $a\in P$ appears in at most one element of $M$.
Definition 2.8.
Let $P$ be a poset, and let $M\subseteq P\times P$ be a partial matching. If there exists no cycle of the following form, then $M$ is called an acyclic matching.
(2.2) $b_{1}\succ a_{1}\prec b_{2}\succ a_{2}\prec\dots\prec b_{p}\succ a_{p}\prec b_{p+1}=b_{1}.$
Here, $(a_{i},b_{i})\in M$ for $i=1,2,\dots,p$, with $a_{i}\neq a_{j}$ and $b_{i}\neq b_{j}$ for $i\neq j$, and $p\geq 2$.
We write $a\longleftrightarrow b$ to denote that two elements $a$ and $b$ of $P$ are matched.
@@
Let $\mathcal{H}(K)$ be the Hasse diagram of $K$, defined as a directed graph whose vertices correspond to the simplices of $K$, and which has a directed edge from $\sigma$ to $\tau$ whenever $\tau<\sigma$ and there is no simplex $\nu$ such that $\tau<\nu<\sigma$. Then, the Morse complex $\mathcal{M}(K)$ of $K$ can be defined as a simplicial complex whose vertices are in one-to-one correspondence to the edges of $\mathcal{H}(K)$, and $n$-simplices correspond to acyclic matching in $\mathcal{H}(K)$ consisting of $n+1$ edges.

### v2

@@ 2. Preliminaries
We write $V(K)\coloneqq\{\sigma^{(0)}\in K\}$ and $E(K)\coloneqq\{\sigma^{(1)}\in K\}$, and refer to them as the vertex set and edge set of $K$, respectively. A 0-simplex adjacent to $v\in V(K)$ is called a neighbor of $v$. The degree of a vertex $v\in V(K)$, denoted $\deg(v)$ is the number of neighbors of $v$. If $\deg(v)=1$ and the neighbor of $v$ is $w$, we say that $\{v,vw\}$ is a leaf.
Definition 2.2.
A matching in $K$ is a subset $M\subset E(K)$ of the edge set such that no two edges in $M$ share a common vertex.
@@
Let $\mathcal{H}(K)$ be the Hasse diagram of $K$, defined as a directed graph whose vertices correspond to the simplices of $K$, and which has a directed edge from $\sigma$ to $\tau$ whenever $\tau<\sigma$ and there is no simplex $\nu$ such that $\tau<\nu<\sigma$. From definition 2.5, we see that the non-critical pairs form a matching in the Hasse diagram. If we reverse the orientation of the arrows in this matching, it can be shown that the resulting directed graph obtained is acyclic. We will call such a matching in the Hasse diagram of the given simplicial complex acyclic matching.
We write $\sigma\longleftrightarrow\tau$ to denote that the simplices $\sigma$ and $\tau$ are matched.
@@
By using Hasse diagrams, we can give an eqivalent definition of the Morse complex as in Definition 2.7: The Morse complex $\mathcal{M}(K)$ of $K$ can be defined as a simplicial complex whose vertices are in one-to-one correspondence to the edges of $\mathcal{H}(K)$, and $n$-simplices correspond to acyclic matching in $\mathcal{H}(K)$ consisting of $n+1$ edges.